solve the quadratic equation by factoring: 15x^2-2x-8=0
Answer by ewatrrr(24785)


Hi,
solving the quadratic equation by factoring:
15x^2-2x-8=0
(5x -4)(3x +2)= 0 Note:SUM of the inner product(-12x) and the outer product(10x) = -2x
(5x -4)(3x +2)= 0
(5x -4)= 0
x = 4/5
(3x +2)= 0
x = -2/3